# Pareto Envelope Augmented with Reinforcement Learning: Multi-objective reinforcement learning-based approach for

### 19 Dec 2023 v2

@@ 5.1.1 Unconstrained test suites
As described in Section 4.2.1, we start by evaluating the performance of the derived algorithms on five classical benchmarks namely dtlz2, dtlz4, dtlz5, dtzl6, and dtlz7. The envelope-based approach PEARL-e described in algorithm 1 is characterized by the generation of different rays every n_steps. Therefore, we must illustrate the role of the rays and the shape of the Dirichlet distribution (by varying values of $\alpha$) on the optimal set of solutions found as well. We also conducted performance comparisons with two legacy GA-based approaches, specifically NSGA-II and NSGA-III. It is worth noting that the computational overhead associated with training and subsequently applying the neural network during inference is likely to be overshadowed by the computational demands of the physics code used in real-world systems, as we observed in our PWR application. Consequently, we have not addressed the latter aspect in this paper.
@@
Table 2: Different HV and associated standard deviation $\sigma$ over 20 experiments using different versions of PEARL against GA-based legacy approaches. For PEARL-e we show different combinations of $\alpha$. The best cases are highlighted in red. Friedman statistical test is significant with p <<much-less-than$<$ < 0.01.
Algorithms dtlz2 dtlz4 dtlz5 dtlz6 dtlz7
PEARL-e $\lambda=0$ 25.96/0.16 25.81/0.14 23.31/0.22 21.86/4.31 33.47/0.11
PEARL-e $\alpha=[0.5,0.5,0.5]$ 26.36/0.03 26.16/0.08 23.86/0.03 24.00/0.001 33.27/0.21
PEARL-e $\alpha=[1,1,1]^{\star}$ 26.40/0.01 26.24/0.06 23.91/0.02 23.99/0.02 33.34/0.32
PEARL-e $\alpha=[10,10,10]$ 26.38/0.01 26.31/0.05 23.87/0.02 23.99/0.001 33.17/0.20
PEARL-e $\alpha=[0.99,2.0,2.0]$ 26.40/0.01 26.26/0.08 23.92/0.01 23.97/0.12 33.09/0.26
PEARL-$\epsilon$ 26.34/0.04 26.16/0.05 23.86/0.04 23.76/0.50 33.35/0.18
PEARL-NdS (crowding) 26.36/0.04 26.15/0.07 23.86/0.04 24.00/0.01 33.15/0.32
NSGA-II 26.340/0.01 26.34/0.01 23.96/0.004 23.97/0.003 32.31/1.75
NSGA-III 26.37/0.01 26.36/0.01 23.93/0.02 23.93/0.01 32.58/0.19
@@
Table 3: Different $I_{C}$ and associated standard deviation $\sigma$ over 20 experiments using different versions of PEARL against GA-based legacy approaches. For PEARL-e we show different combinations of $\alpha$. The best cases are highlighted in red. Friedman statistical test is significant with p <<much-less-than$<$ < 0.01.
Algorithms dtlz2 dtlz4 dtlz5 dtlz6 dtlz7
PEARL-e $\lambda=0$ 0.05/0.22 0.0/0.0 33.75/7.25 1851.5/1780.1 5217.8/1122.9
PEARL-e $\alpha=[0.5,0.5,0.5]$ 542.65/265.60 2.2/2.99 35.1/4.59 12043.0/1273.8 4434.8/1593.7
PEARL-e $\alpha=[1,1,1]$ 1478.5/253.20 11.55/8.26 38.05/8.35 13602/1807.0 4240.6/1223.7
PEARL-e $\alpha=[10,10,10]$ 3359.0/828.08 32.7/16.50 371.6 /219.57 14362.0/1027.1 2293.7/1232.7
PEARL-e $\alpha=[0.99,2.0,2.0]$ 1908.8/279.30 14.1/9.22 51.05/14.04 13375./1115.1 3697.4/1262.1
PEARL-$\epsilon$ 478.45/238.1 2 1.5/1.12 35.25/4.58 10938./3643.5 3976.9/1187.9
PEARL-NdS (crowding) 503.35/246.28 2.05/2.52 34.8/5.80 11468./2449.8 4145.7/1154.8
NSGA-II 46.65/3.86 46.3/1.85 48.1/3.97 57.3/0.78 10.75/6.92
NSGA-III 43.6/2.33 34.3/3.82 43.55/3.89 51.4/2.44 16.25/5.85
@@
Additionally, you can find associated Figures depicting the optimal Pareto front and the solutions obtained by various algorithms in appendix A.1 and in particular Figures 6 and 7. There, we examined the visual impact of $\alpha$ on Figure 6, demonstrating the superiority of PEARL-e if the location of the points in the pareto front is known, but highlighting the drawback of requiring this prior knowledge. On the other hand, PEARL-NdS consistently performs well without relying on assumptions about the studied optimum surface. This provides valuable insights into the algorithm’s robustness. This plays a major role for real-world application in which we do not usually have knowledge of how the front looks like.
Lastly, we can analyze the NT tests results for the HVs in Table 4. For the sake of brevity, we’ve only included the columns corresponding to the algorithm with the highest average HV across the 20 experiments. For each case where PEARL is superior (dtlz2, dtlz6 and dtlz7), the NTs are successful against GA-based legacy MOO methods. On the other hand, for problems for which NSGA-based approaches are superior (dtlz4 and dtlz5), some NTs failed against PEARL (e.g., PEARL-e with $\alpha$ $=[0.99,2.0,2.0]$ for dtlz5). This implies that there are no statistically significant differences between these algorithms on the latter problems but for the former problems PEARL is statistically better. This further underscores the potential for PEARL to serve as a viable alternative to legacy approaches in unconstrained MOO scenarios.
@@
dtlz2 PEARL-e $\alpha=[1,1,1]$ dtlz4 NSGA-III dtlz5 NSGA-II dtlz6 PEARL-e $\alpha=[0.5,0.5,0.5]$ dtlz7 PEARL-e $\lambda=0$
PEARL-e $\lambda=0$ 0.001 0.001 0.001 0.001 -
PEARL-e $\alpha=[0.5,0.5,0.5]$ 0.001 0.001 0.001 - 0.212
PEARL-e $\alpha=[1,1,1]$ 0.001 0.034 0.9 0.819
PEARL-e $\alpha=[10,10,10]$ 0.264 0.324 0.001 0.086 0.005
PEARL-e $\alpha=[0.99,2.0,2.0]$ 0.9 0.005 0.357 0.626 0.001
PEARL-NdS (crowding) 0.002 0.001 0.001 0.357 0.007
PEARL-$\epsilon$ 0.001 0.001 0.001 0.392 0.691
NSGA-II 0.001 0.9 - 0.001 0.001
NSGA-III 0.002 - 0.8192 0.001 0.001
@@ 5.1.2 Constrained test suites
We conduct similar experiments as in Section 5.1.1, where we provide the HV and $I_{C}$. We are using 10,000 steps for all problems. The input space $\mathcal{A}$ is such that $\mathcal{A}=[0,1]^{n_{x}}\in\mathbb{R}$ where $n_{x}$ is equal to 7 for the cxdtlzy, while all ctpx have an input dimension of 2. The nadir is arbitrarily chosen to be $[3,3,3]$ and $[3,3]$ for the cxdtlzy and ctpx, respectively. ncores and normalized_obj are set to 8 and False, respectively. We use uniformity as cosine similarity cos and $\lambda$ = 1. $\alpha$ $=[10,10,10]$, $\kappa$ = 64 for all but niching2 for which we choose 1024, M equal to $\kappa$, and $\nu$ = 0.05. We compare the approaches to the constrained version of NSGA-III. The results are given in Tables 5 and 6. For statistical rigour, we provide the average and standard deviation $\sigma$ over 20 experiments for each.
@@
Looking at the higher HV and $I_{C}$ for most CL-based PEARL, we see that utilizing the CL method (i.e., solving for the constraints first) shows great performance for all but the c2dtlz2 problem with the PEARL-e method, while NSGA performs well throughout. NSGA-III [17] is also treating constraints first before dealing with the regular target FOMs. niching2, which is a CL-based approach replacing the distance to the feasible space by a rank (see algorithm 2),and PEARL-NdS (crowding) present the highest HV for 3 out of the 6 benchmarks and the highest $I_{C}$ for one. The hypothesis put forth in Section 3.4, suggesting to divide the problem into two simpler sub-problems, has indeed led to high performance for the RL agents. Furthermore, across all benchmark scenarios, the HV and $I_{C}$ are greater than NSGA-III for at least one version of PEARL (with CL). It confirms therefore the potential of PEARL (with CL) to scale for different problems with different front distributions in the constrained case as well.
The associated figure depicting the optimum Pareto front and the solution found by the various algorithms are given in appendix A.2 and in particular Figure 8. There, we see that even if the number of Pareto-optimal solutions is small (see Table 6) on problems with very difficult pareto front (e.g., ctpx problems), many feasible points close to them could be found with PEARL compared to the NSGA-III approach.
@@
Table 5: Comparison of HV and associated standard deviation $\sigma$ over 20 experiments using different versions of C-PEARL against GA-based legacy approaches. for the application to classical benchmarks. The best cases are highlighted in red. Friedman statistical test is significant with p <<much-less-than$<$ < 0.01.
algorithm c2dtlz2 c3dtlz4 CTP1 CTP2 CTP3 CTP4
PEARL-e 26.13 / 0.03 23.23 / 0.23 7.20 / 0.01 7.71 / 0.01 7.61 / 0.01 7.04/ 0.22
PEARL-$\epsilon$ 22.65 / 2.56 23.09 / 0.26 6.21 / 0.06 7.76 / 0.004 7.66 / 0.01 7.18 / 0.23
PEARL-NdS (crowding)-CL 26.30 / 0.02 23.10 / 0.21 7.22 / 0.01 7.75 / 0.01 7.64 / 0.01 7.03 / 0.18
PEARL-NdS (niching2) 26.15/0.02 23.56/0.21 7.22/0.01 7.74/0.01 7.63/0.01 7.28/0.15
NSGA-III 25.07 / 1.29 22.97 / 1.83 6.96/0.05 7.73/0.01 7.64 /0.02 6.63/0.53
@@
Table 6: Comparison of $I_{C}$ and associated standard deviation $\sigma$ over 20 experiments using different versions of C-PEARL against GA-based legacy approaches for the application to classical benchmarks. The best cases are highlighted in red. All FTs succeeded with p = 0.00 except for ctp3 and ctp4.
algorithm c2dtlz2 c3dtlz4 CTP1 CTP2 CTP3 CTP4
PEARL-e 10.6 / 7.85 12.25 / 4.59 45.45 / 49.77 1.35/ 0.48 1 / 0 1/ 0
PEARL-$\epsilon$ 39.75 / 49.46 8 / 2.63 520 / 0 9.8 / 4.73 1.05/ 0.2 1 /0
PEARL-NdS (crowding)-CL 192.1 / 59.84 12.05 /4.33 321.7/ 19.58 18.15 /6.20 1.05/ 0.22 1/ 0
PEARL-NdS (niching2) 12.0/4.40 19.4/4.57 83.35/12.86 2.6/1.53 1 /0.0 1 /0.0
NSGA-III 34.15 / 7.21 37.20 / 13.97 8.55 / 1.47 2.45 / 1.24 1.05 / 0.21 0.95/ 0.22
@@
Finally, we can analyze the NTs results for the HV in Table 7. To streamline the presentation, we’ve focused on the columns corresponding to the algorithm with the highest average HV across the 20 experiments in each case. In every scenario, one version of PEARL consistently outperforms others by achieving the highest average HV.
@@
Furthermore, in 4 out of the 6 benchmarks, the NTs demonstrate that PEARL surpasses the NSGA-III approach. This underscores the potential for PEARL to serve as a successful alternative to legacy approaches in constrained MOO scenarios. Notably, the results are even more favorable in the constrained cases compared to the unconstrained ones (as shown in appendix 5.1.1).
@@
c2dtlz2 (crowding) with CL c3dtlz4 (niching2) ctp1 (crowding) with CL ctp2 PEARL-$\epsilon$ with CL ctp3 PEARL-$\epsilon$ with CL ctp4 (niching2)
PEARL-e ($\alpha=[10,10,10]$) 0.00290 0.530 0.147 0.003 0.001 0.039
PEARL-NdS (crowding)-CL - 0.464 - 0.9 0.883 0.530
PEARL-NdS (niching)-CL 0.9 0.264 0.9 0.9 0.787 0.147
PEARL-NdS (niching2) 0.00357 - 0.755 0.357 0.020 -
PEARL-$\epsilon$ 0.001 0.129 0.001 - - 0.9
NSGA-III 0.001 0.9 0.001 0.046 0.188 0.001

### 15 Mar 2024 v3

@@ 5.1.1 Unconstrained test suites
As described in Section 4.2.1, we start by evaluating the performance of the derived algorithms on seven classical benchmarks namely dtlz1, dtlz2, dtlz3, dtlz4, dtlz5, dtzl6, and dtlz7. The envelope-based approach PEARL-e described in algorithm 1 is characterized by the generation of different rays every n_steps. Therefore, we must illustrate the role of the rays and the shape of the Dirichlet distribution (by varying values of $\alpha$) on the optimal set of solutions found as well. We also conducted performance comparisons with four legacy approaches, specifically NSGA-II, NSGA-III, SA, and TS. It is worth noting that the computational overhead associated with training and subsequently applying the neural network during inference is likely to be overshadowed by the computational demands of the physics code used in real-world systems, as we observed in our single-objective PWR application [53]. Consequently, we have not addressed the latter aspect in this paper.
@@
Table 2: Different HV and associated standard deviation $\sigma$ over 20 experiments using different versions of PEARL against SO-based legacy approaches. For PEARL-e we show different combinations of $\alpha$. The best cases are highlighted in red. Friedman statistical test is significant with p <<much-less-than$<$ < 0.05.
Algorithms dtlz1 dtlz2 dtlz3 dtlz4
PEARL-e $\lambda=0$ 26.72/ 0.14 25.96/0.16 23.53/ 3.04 25.81/0.14
PEARL-e $\alpha=[0.5,0.5,0.5]$ 26.82/ 0.11 26.36/0.03 24.68/ 1.44 26.16/0.08
PEARL-e $\alpha=[1,1,1]^{\star}$ 26.71/ 0.17 26.40/0.01 23.15/ 1.93 26.24/0.06
PEARL-e $\alpha=[10,10,10]$ 26.49/ 0.41 26.38/0.01 21.35/ 5.37 26.31/0.05
PEARL-e $\alpha=[0.99,2.0,2.0]$ 26.73 / 0.12 26.40/0.01 23.90 / 2.28 26.26/0.08
PEARL-$\epsilon$ 26.93/ 0.02 26.34/0.04 26.00/ 0.14 26.16/0.05
PEARL-NdS (niching) 26.93/ 0.01 26.33 / 0.01 25.98/ 0.16 26.20 / 0.03
PEARL-NdS (crowding) 26.93/ 0.02 26.36/0.04 26.02/ 0.14 26.15/0.07
NSGA-II 26.96/ 0.04 26.340/0.01 26.36/ 0.01 26.35/ 0.01
NSGA-III 26.96/ 0.04 26.37/0.01 26.37/ 0.01 26.36/0.01
SA 26.69/ 0.27 26.04 / 0.07 23.07/ 4.39 25.07 / 0.65
TS 26.88/ 0.06 26.03 / 0.09 25.67/ 0.65 23.82 / 0.98
@@
Table 3: Different HV and associated standard deviation $\sigma$ over 20 experiments using different versions of PEARL against SO-based legacy approaches. For PEARL-e we show different combinations of $\alpha$. The best cases are highlighted in red. Friedman statistical test is significant with p <<much-less-than$<$ < 0.05.
Algorithms dtlz5 dtlz6 dtlz7
PEARL-e $\lambda=0$ 23.31/0.22 21.86/4.31 33.47/0.11
PEARL-e $\alpha=[0.5,0.5,0.5]$ 23.86/0.03 24.00/0.001 33.27/0.21
PEARL-e $\alpha=[1,1,1]$ 23.91/0.02 23.99/0.02 33.34/0.32
PEARL-e $\alpha=[10,10,10]$ 23.87/0.02 23.99/0.001 33.17/0.20
PEARL-e $\alpha=[0.99,2.0,2.0]$ 23.92/0.01 23.97/0.12 33.09/0.26
PEARL-$\epsilon$ 23.86/0.04 23.76/0.50 33.35/0.18
PEARL-NdS (niching) 23.79/0.03 23.99/0.00 33.37 / 0.08
PEARL-NdS (crowding) 23.86/0.04 24.00/0.01 33.15/0.32
NSGA-II 23.96/0.004 23.97/0.003 32.31/1.75
NSGA-III 23.93/0.02 23.93/0.01 32.58/0.19
SA 23.73  0.06 -289.85 / 12.09 20.59 / 1.82
TS 23.74  0.16 -92.02 / 34.46 30.17 / 0.55
Table 4: Different $I_{C}$ and associated standard deviation $\sigma$ over 20 experiments using different versions of PEARL against SO-based legacy approaches. For PEARL-e we show different combinations of $\alpha$. The best cases are highlighted in red. Friedman statistical test is significant with p <<much-less-than$<$ < 0.05.
Algorithms dtlz1 dtlz2 dtlz3 dtlz4
PEARL-e $\lambda=0$ 1.4/ 1.2 0.05/0.0 0.9 / 0.94 0.0 / 0.0
PEARL-e $\alpha=[0.5,0.5,0.5]$ 1.2 /0.93 542.65/265.60 1.25 / 0.99 2.2/2.99
PEARL-e $\alpha=[1,1,1]$ 0.8 /0.75 1478.5/253.20 0.4 /0.9 11.55/8.26
PEARL-e $\alpha=[10,10,10]$ 0.85/ 0.79 3359.0/828.08 0.6 / 0.8 32.7/16.50
PEARL-e $\alpha=[0.99,2.0,2.0]$ 1.3/1.38 1908.8/279.30 0.7/ 0.78 14.1/9.22
PEARL-$\epsilon$ 3.95 / 1.72 478.45/238.1 2.05/ 1.40 2 1.5/1.12
PEARL-NdS (niching) 4.6 / 2.27 26.55/ 9.02 2.35 / 1.28 2.05/2.13
PEARL-NdS (crowding) 5.7 /3.03 503.35/246.28 2.35/ 1.8 2.05/2.52
NSGA-II 51.8/ 11.91 46.65/3.86 53.25/ 0.8 46.3/1.85
NSGA-III 49.9/ 11.71 43.6/2.33 52 /2.81 34.3/3.82
SA 36.5/ 51.89 11.95 / 6.52 24.4/ 48.11 14.6 / 3.64
TS 79.1/ 53.63 47.65 / 13.39 45.4/ 53.83 5.35 / 2.15
Table 5: Different $I_{C}$ and associated standard deviation $\sigma$ over 20 experiments using different versions of PEARL against SO-based legacy approaches. For PEARL-e we show different combinations of $\alpha$. The best cases are highlighted in red. Friedman statistical test is significant with p <<much-less-than$<$ < 0.05.
Algorithms dtlz5 dtlz6 dtlz7
PEARL-e $\lambda=0$ 33.75/7.25 1851.5/1780.1 5217.8/1122.9
PEARL-e $\alpha=[0.5,0.5,0.5]$ 35.1/4.59 12043.0/1273.8 4434.8/1593.7
PEARL-e $\alpha=[1,1,1]$ 38.05/8.35 13602/1807.0 4240.6/1223.7
PEARL-e $\alpha=[10,10,10]$ 371.6 /219.57 14362.0/1027.1 2293.7/1232.7
PEARL-e $\alpha=[0.99,2.0,2.0]$ 51.05/14.04 13375./1115.1 3697.4/1262.1
PEARL-$\epsilon$ 35.25/4.58 10938./3643.5 3976.9/1187.9
PEARL-NdS (niching) 28.35/6.01 499.1/1.3 352.85/13.87
PEARL-NdS (crowding) 34.8/5.80 11468./2449.8 4145.7/1154.8
NSGA-II 48.1/3.97 57.3/0.78 10.75/6.92
NSGA-III 43.55/3.89 51.4/2.44 16.25/5.85
SA 0.1 / 0.44 0 / 0 0 / 0
TS 19.15 / 8.18 0 / 0 0/0
@@
Additionally, while PEARL-e proves to be an efficient method (high HV on 6 out of 7 unconstrained benchmarks), its performance depends on the choice of $\alpha$, hence we must assume some knowledge of the location of the Pareto front. This accentuates the challenge of potentially introducing computational overhead to adapt the rays to the specific problem at hand. This complexity may arise, particularly when the Pareto front is unknown. However, when a designer possesses some knowledge of the Pareto surface, fine-tuning the Dirichlet distribution with $\alpha$ could yield significant advantages. Conversely, the PEARL-NdS methods systematically found Pareto-optimal solutions without the need for any assumptions regarding ray distributions tailored to a specific front. This flexibility is highly beneficial for real-world practitioners who seek to utilize these tools without delving into problem-specific tuning.
Lastly, we can analyze the NTs results for the HVs in Tables 6 and 7. For the sake of brevity, we’ve only included the columns corresponding to the algorithm with the highest average HV across the 20 experiments. For each case where PEARL is superior (dtlz2, dtlz6 and dtlz7), the NTs are successful against SO-based legacy MOO methods. On the other hand, for problems for which NSGA-based approaches are superior (dtlz1, dtlz3, dtlz4, and dtlz5), some NTs failed against PEARL (e.g., PEARL-e with $\alpha$ $=[0.99,2.0,2.0]$ for dtlz5). This implies that there are no statistically significant differences between these algorithms on the latter problems but for the former problems PEARL is statistically better. This further underscores the potential for PEARL to serve as a viable alternative to legacy approaches in unconstrained MOO scenarios.
@@
dtlz1 NSGA-II dtlz2 PEARL-e $\alpha=[1,1,1]$ dtlz3 NSGA-III dtlz4 NSGA-III
PEARL-e $\lambda=0$ 0.001 0.001 0.001 0.001
PEARL-e $\alpha=[0.5,0.5,0.5]$ 0.001 0.001 0.001 0.001
PEARL-e $\alpha=[1,1,1]$ 0.001 - 0.001 0.001
PEARL-e $\alpha=[10,10,10]$ 0.001 0.264 0.001 0.324
PEARL-e $\alpha=[0.99,2.0,2.0]$ 0.001 0.9 0.001 0.005
PEARL-$\epsilon$ 0.9 0.001 0.053 0.001
PEARL-NdS (niching) 0.9 0.001 0.16 0.054
PEARL-NdS (crowding) 0.68 0.002 0.053
NSGA-II - 0.001 0.9 0.1
NSGA-III 0.9 0.002 - -
SA 0.001 0.001 0.001 0.001
TS 0.02 0.001 0.004 0.003
dtlz5 NSGA-II dtlz6 PEARL-e $\alpha=[0.5,0.5,0.5]$ dtlz7 PEARL-e $\lambda=0$
PEARL-e $\lambda=0$ 0.001 -
PEARL-e $\alpha=[0.5,0.5,0.5]$ 0.001 - 0.212
PEARL-e $\alpha=[1,1,1]$ 0.034 0.9 0.819
PEARL-e $\alpha=[10,10,10]$ 0.001 0.9 0.04
PEARL-e $\alpha=[0.99,2.0,2.0]$ 0.357 0.626 0.001
PEARL-$\epsilon$ 0.001 0.392 0.691
PEARL-NdS (crowding) 0.001 0.357 0.007
PEARL-NdS (niching) 0.001 0.001 0.9
NSGA-II - 0.001 0.001
NSGA-III 0.8192 0.001 0.001
SA 0.001 0.001 0.001
TS 0.001 0.001 0.001
@@ 5.1.2 Constrained test suites
We conduct similar experiments as in Section 5.1.1, where we provide the HV and $I_{C}$. We are using 10,000 steps for all problems. The input space $\mathcal{A}$ is such that $\mathcal{A}=[0,1]^{n_{x}}\in\mathbb{R}$. The nadir is arbitrarily chosen to be $[3,3,3]$ and $[3,3]$ for the cxdtlzy and ctpx, respectively. ncores and normalized_obj are set to 8 and False, respectively. We use uniformity as cosine similarity cos and $\lambda$ = 1. $\alpha$ $=[10,10,10]$, $\kappa$ = 64 for all but niching2 for which we choose 1024, M equal to $\kappa$, and $\nu$ = 0.05. We compare the approaches to the constrained version of NSGA-III, SA, and TS as described in Section 4.2.3. The results are given in Tables 8, 9, 10, and 11. For statistical rigour, we provide the average and standard deviation $\sigma$ over 20 experiments for each.
@@
Looking at the higher HV and $I_{C}$ for most CL-based PEARL, we see that utilizing the CL method (i.e., solving for the constraints first) shows great performance for all but the c2dtlz2 problem with the PEARL-e method. The hypothesis put forth in Section 3.4, suggesting to divide the problem into two simpler sub-problems, has indeed led to high performance for the RL agents. NSGA performs well throughout except for c1dtlz1 and c1dtlz3. NSGA-III [17] is also treating constraints first before dealing with the regular target FOMs. PEARL-NdS (crowding) and (crowding) present the highest HV for 4 out of the 8 benchmarks and the former the highest $I_{C}$ for 5 out of 8. Furthermore, across all benchmark scenarios, the HV and $I_{C}$ are greater than SO-based techniques for at least one version of PEARL (with CL). It confirms therefore the potential of PEARL (with CL) to scale for different problems with different front distributions in the constrained case as well.
@@
Table 8: Comparison of HV and associated standard deviation $\sigma$ over 20 experiments using different versions of C-PEARL against SO-based legacy approaches. for the application to classical benchmarks. The best cases are highlighted in red. Friedman statistical test is significant with p <<much-less-than$<$ < 0.05.
algorithm c1dtlz1 c1dtlz3 c2dtlz2 c3dtlz4
PEARL-e 17.09/ 9.88 18.48 7.76 26.13 / 0.03 23.23 / 0.23
PEARL-$\epsilon$ 23.53/ 5.50 25.97 / 0.28 22.65 / 2.56 23.09 / 0.26
PEARL-NdS (niching)-CL 23.64/ 5.51 25.84 0.26 26.31 /0.02 23.19 /0.23
PEARL-NdS (crowding)-CL 24.45/ 1.32 26.03 / 0.18 26.30 / 0.02 23.10 / 0.21
PEARL-NdS (niching2) 24.21/ 1.00 17.99 0.05 26.15/0.02 23.56/0.21
PEARL-NdS (crowding2) 24.97/ 0.88 17.97  0.18 18.00 /0.03 23.56 /0.19
NSGA-III 3.37/ 8.03 17.60 0.96 25.07 / 1.29 22.97 / 1.83
SA 17.80/ 8.90 12.43 5.40 17.86 / 0.08 14.42 / 2.20
TS 11.18/ 11.19 15.23 3.51 18.00 / 0.00 14.73 / 2.45
@@
Table 9: Comparison of HV and associated standard deviation $\sigma$ over 20 experiments using different versions of C-PEARL against SO-based legacy approaches. for the application to classical benchmarks. The best cases are highlighted in red. Friedman statistical test is significant with p <<much-less-than$<$ < 0.05.
algorithm ctp1 ctp2 ctp3 ctp4
PEARL-e 7.20 / 0.01 7.71 / 0.01 7.61 / 0.01 7.04/ 0.22
PEARL-$\epsilon$ 6.21 / 0.06 7.76 / 0.004 7.66 / 0.01 7.18 / 0.23
PEARL-NdS (niching)-CL 7.21 /0.02 7.76 /0.00 7.65 /0.01 7.11 /0.21
PEARL-NdS (crowding)-CL 7.22 / 0.01 7.75 / 0.01 7.64 / 0.01 7.03 / 0.18
PEARL-NdS (niching2) 7.22/0.01 7.74/0.01 7.63/0.01 7.28/0.15
PEARL-NdS (crowding2) 7.22 /0.00 7.76 /0.00 7.67 /0.01 7.33 /0.18
NSGA-III 6.96/0.05 7.73/0.01 7.64 /0.02 6.63/0.53
SA 7.16 / 0.02 7.73 / 0.01 7.58 / 0.07 7.19 / 0.12
TS 7.19 / 0.01 7.76/ 0.01 7.63 / 0.04 7.28 / 0.07
Table 10: Comparison of $I_{C}$ and associated standard deviation $\sigma$ over 20 experiments using different versions of C-PEARL against SO-based legacy approaches for the application to classical benchmarks. The best cases are highlighted in red. All FTs succeeded with p = 0.00.
algorithm c1dtlz1 c1dtlz3 c2dtlz2 c3dtlz4
PEARL-e 0.45/ 0.80 0.35/ 0.65 10.6 / 7.85 12.25 / 4.59
PEARL-$\epsilon$ 2.15/ 1.96 3.70 /3.07 39.75 / 49.46 8 / 2.63
PEARL-NdS (niching)-CL 2.25/ 1.22 2.40 /1.46 184.65 /45.48 11.85/ 3.07
PEARL-NdS (crowding)-CL 2.55/ 1.99 2.95/ 1.69 192.1 / 59.84 12.05 /4.33
PEARL-NdS (niching2) 2.70/ 1.79 0.00/ 0 12.0/4.40 19.4/4.57
PEARL-NdS (crowding2) 4.25/ 1.44 0.00/ 0 0 /0 17.15/ 3.90
NSGA-III 0.00/ 0.00 0.00/ 0 34.15 / 7.21 37.20 / 13.97
SA 0.45/ 1.75 0.20 / 0.68 0 / 0 10 / 3.11
TS 0.00/ 0.00 0.00/ 0 0.8 / 2.66 10.45/ 5.65
Table 11: Comparison of $I_{C}$ and associated standard deviation $\sigma$ over 20 experiments using different versions of C-PEARL against SO-based legacy approaches for the application to classical benchmarks. The best cases are highlighted in red. All FTs succeeded with p = 0.00.
algorithm ctp1 ctp2 ctp3 ctp4
PEARL-e 45.45 / 49.77 1.35/ 0.48 1 / 0 1/ 0
PEARL-$\epsilon$ 520 / 0 9.8 / 4.73 1.05/ 0.2 1 /0
PEARL-NdS (niching)-CL 344.8/ 49.22 14.7 /4.86 1.05 /0.22 1.05/ 0.22
PEARL-NdS (crowding)-CL 321.7/ 19.58 18.15 /6.20 1.05/ 0.22 1/ 0
PEARL-NdS (niching2) 83.35/12.86 2.6/1.53 1 /0.0 1 /0.0
PEARL-NdS (crowding2) 366.95/ 13.99 13.4 /2.97 1 /0 1 /0
NSGA-III 8.55 / 1.47 2.45 / 1.24 1.05 / 0.21 0.95/ 0.22
SA 11.4 / 8.73 2.25 / 1.813 0.00 / 0.00 0.05 / 0.22
TS 26.75 / 9.65 6.2 / 2.91 0.00 / 0.00 0.05 / 0.22
@@
Finally, we can analyze the NTs results for the HV in Tables 12 and 13. To streamline the presentation, we’ve focused on the columns corresponding to the algorithm with the highest average HV across the 20 experiments in each case. In every scenario, one version of PEARL consistently outperforms others by achieving the highest average HV.
@@
Furthermore, in at least 6 out of the 8 benchmarks, the NTs demonstrate that PEARL surpasses the SO-based approaches. This underscores the potential for PEARL to serve as a successful alternative to legacy approaches in constrained MOO scenarios. Notably, the results are even more favorable in the constrained cases compared to the unconstrained ones.
@@
c1dtlz1 (crowding2) c1dtlz3 (crowding) with CL c2dtlz2 (crowding) with CL c3dtlz4 (niching2)
PEARL-e ($\alpha=[10,10,10]$) 0.001 0.00 0.00290 0.530
PEARL-$\epsilon$ 0.9 0.9 0.001 0.129
PEARL-NdS (niching)-CL 0.9 0.9 0.9 0.264
PEARL-NdS (crowding)-CL 0.9 - - 0.11
PEARL-NdS (niching2) 0.9 0.001 0.00357 -
PEARL-NdS (crowding2) - 0.001 0.001 0.9
NSGA-III 0.001 0.001 0.001 0.9
SA 0.001 0.001 0.001 0.001
TS 0.001 0.001 0.001 0.001
ctp1 (crowding) with CL ctp2 PEARL-$\epsilon$ with CL ctp3 PEARL-$\epsilon$ with CL ctp4 (niching2)
PEARL-e ($\alpha=[10,10,10]$) 0.147 0.003 0.001 0.039
PEARL-$\epsilon$ 0.001 - - 0.9
PEARL-NdS (niching)-CL 0.9 0.9 0.787 0.147
PEARL-NdS (crowding)-CL - 0.9 0.12 0.06
PEARL-NdS (niching2) 0.755 0.357 0.020 -
PEARL-NdS (crowding2) 0.84 0.9 0.9 0.9
NSGA-III 0.001 0.046 0.188 0.001
SA 0.001 0.001 0.001 0.9
TS 0.001 0.9 0.00 0.9